# On the Welfare of EIP-1559 with Patient Bidders

### v1

@@ 3.3 Proof of The Main Lemma
Since $\beta>1$ is the first time the price is at least $\theta^{\prime}$, the price was rising from time $\beta-1$, and $p_{\beta-1}<\theta^{\prime}=\theta\cdot e^{-\eta}$. In this case $p_{\beta}$ is at most $\theta^{\prime}\cdot e^{\eta\cdot(c-1)}$. As once the price get to $\theta^{\prime}$ it never drops below it, we have $p_{T+\Gamma+1}\geq\theta^{\prime}$. We conclude that $\frac{p_{\beta}}{p_{T+\Gamma+1}}\leq\frac{\theta^{\prime}\cdot e^{\eta\cdot(c-1)}}{\theta^{\prime}}=e^{\eta\cdot(c-1)}$ and therefore
@@ 4 Lower Bounds
Our lower bounds apply to algorithms that, like the EIP-1559 algorithm, are allowed both some slackness $\Delta$ and some extension $\Gamma$ in order to compete with the optimum schedule. As these two notions of augmentation may be traded-off against each other to some extent [^24] All our lower bounds hold even for a single horizon $T$ that is known to the algorithm in advance. In that setting, shifting between the two parameters is indeed possible. If there is a single horizon $T$ that is known by the algorithm, the extension may be completely transferred to additional slackness by immediately scheduling at the last block all transactions that were intended for the extension period. Similarly, slackness can be shift to an extension when $T$ is known, by postponing some scheduling to the extension time., we will generally prove lower bounds on their sum.
@@ 4.1 Augmentation of Price-Based Algorithms must depend on Value Range
We will look at our algorithm running for $R=T+\Gamma$ steps, where all transactions arrive at time $1$. We will have $12^{R}+$ 1 possible inputs scenarios, where in input scenario number $m$ ($0\leq m\leq 2^{R}$) we have, for every integer $0\leq i\leq m$, exactly $(R+\Delta)$ transactions with value $r^{i}$ (all arriving at time 1), each of size $1$. We also specify how our adversary will respond to every price $p$ in each input scenario $m$. If $p>r^{m}$ then there are no transactions that are willing to pay this price so no transactions will be scheduled. if $p\leq r^{m}$ then the adversary will always schedule only transactions with the lowest possible value (i.e. those with value $r^{i}$ for the smallest $i$ so that $r^{i}\geq p$). Note that there are enough transactions of this value to suffice for anything that can be scheduled by the end of time $R$.
Now let us look at a price-based algorithm running for $R$ steps on one of these input scenarios. For any price $p$ that is queried by the algorithm there are at most two possible answers by the adversary on any one of these scenarios: either no transaction is scheduled or as many transactions as allowed by the block’s capacity, each of the value $r^{i}$ for the $i$ completely determined by the price, is scheduled. It follows that the complete set of possible behaviors of the algorithm on all possible scenarios can be described as a binary tree of depth $R$, containing at most $12^{R}-$ 1 possible prices in all and splitting the possible transactions into the $2^{R}$ leaves of the tree. Since there are $12^{R}+$ 1 possible scenarios, two of them must map into the same leaf and thus produce the exact same schedule. So let us assume that input scenarios $m<m^{\prime}$ get the same behavior by the algorithm, which must thus never schedule any transactions with any value strictly greater than $r^{m}$ as such do not exists in scenario $m$.
@@
Now for our choice of parameters. We let $T=\Gamma+\Delta$ which ensures that the nominator in the expression above is the constant $2$, and thus, the fraction of social welfare (with respect to the benchmark of any schedule with worst-case block size limit $1$, and no slackness or augmentation) that is achieve by the algorithm is bounded from above by $2/r$. Our choice of $T$ also ensures that $R=T+\Gamma\leq 2T=2(\Gamma+\Delta)$ and since for our choice of $r$ we only needed that $r^{2^{R}}\leq H$ we may choose $r=H^{1/2^{2T}}=H^{1/2^{2(\Gamma+\Delta)}}=H^{1/4^{(\Gamma+\Delta)}}$, which concludes the proof. ∎
@@ 4.2 Online Algorithms need to Relax the Maximum Block Size
Any online algorithm that produces schedules with maximum block size $c\cdot B$, for some $c<2$ and average block size limit $B$ with slackness $\Delta(T)=o(T)$, even with an extension $\Gamma(T)=o(T)$, must lose at least $\delta_{0}=min(1/8,(2-c)/3)$ fraction of social welfare relative to schedules with worst-case block size limit $B$ (even when the values of all transactions are in {1,2}).
@@
Case I: $G\leq T/4$. In this case the total value achieved in the first $T/2$ steps is at most $42\cdot G+1\cdot(1/2-G)\leq 3/$ 4. In this case consider the case where for the next $T/2$ steps, in each step a new transaction of size 1 and value $2$ enters. In this case the best that the mechanism can do is get value $2$ for each of the remaining $T/2$ steps and the extension period $\Gamma=o(T)$ for a total social welfare of at most $7T/4+o(T)$. In comparison the optimal schedule with block size 1 would schedule all the green transactions in the first half, getting a value of $2$ every step for a total of $2T$, so the ratio of welfares in this case is bounded from above by $7/8+o(1)$.
Case II: $G>T/4$. In this case consider the case where for the next $T/2$ steps we get a large number of small transactions of size $\gamma$ with value $\gamma$ (i.e. value per unit of $1$) each. The total size of all blocks produced by our algorithm during all $T$ steps and the extension $\Gamma$ is at most $T+\Delta+\Gamma=T+o(T)$. The total welfare achieved by the algorithm can thus be bounded by $T+o(T)$ (obtained from red transactions and from small transactions) plus $T/2$ (the extra 1 value from every green transaction) minus $(1-c/2-\varepsilon)\cdot G$ which is the capacity completely lost by scheduling green transactions in the first half. Thus the total welfare is at most $T\cdot(3/2+c/2-1+\varepsilon+o(1))$. The optimal schedule with maximum block size 1 would keep all the green transactions for the second half, getting a social welfare of exactly $23T/$ 2. The ratio between these two welfare expressions is $(1+c+o(1))/3$, so the fraction lost is at least $31-(1+c+o(1))/3=(2-c-o(1))/$ 3. ∎
@@ 4.3.1 $c=2$ Does not Suffice for the EIP-1559 Algorithm
In Proposition 4.2 we showed that the maximum block size must be chosen to be at least $2\cdot B$ in any online algorithm that attempts to compete with offline schedules whose worst case block size is $B$ (when $q_{max}=B$), i.e., in our notation, the lower bound showed that we need $c\geq 2$. Proposition 3.1 showed that $c=2$ indeed suffices for the online algorithm that we presented. However, our main theorem for $q_{max}=B$ requires the maximum block size to be strictly greater than twice the average block size limit, $c>2$, and in fact the result also uses an extension that grows in rate that is inversely proportional to $c-2$. So the question is whether having the maximum block size $2\cdot B$ in the EIP-1559 algorithm, i.e. $c=2$, suffices for near-optimally. This question is especially interesting since the actual choice of parameters in Ethereum is exactly $c=2$, as the target block size is $B=15M$ gas and the maximum block size is defined to be $2B=30M$ gas.
@@
Here is an example showing that the EIP-1559 algorithm with $c=2$, even with an extension that grows as $\Gamma(T)=o(T)$, cannot give even half of the optimum social welfare.252525 This example also shows that even when $c>2$, as $c$ approaches $2$, the extension length must grow inversely proportional to $c-2$. Simple tweaks in the example will show that even longer extensions cannot guarantee even a lower fraction of the welfare. For notional convenience let us assume without loss of generality that $p_{min}=1$ (otherwise just scale all the values in the rest of the proof by $p_{min}$). Our example starts with steps where no transactions arrive, and so blocks are empty and the price decreases to $p_{min}=1$. From this point on, every time step $t$, we get as input one “high” transaction of size exactly $B$ and value per unit $10$, and additional “low” transactions of total size $(1+\varepsilon)\cdot B$ and value per unit $2$, where $\varepsilon>0$ is a parameter, to be specified below, that may be as small as we wish. (The size composition of the low transactions does not matter as long as their total size is $(1+\varepsilon)\cdot B$.) In our example the low transactions offer a high tip, i.e. the adversary always schedules them before the high transaction.
@@
Now let us see how the EIP-1559 algorithm deals with this input. As long as $p_{t}\leq 2$, both the high and the low transactions can be scheduled, and since the low transactions have priority they will be scheduled first, filling the block to size $(1+\varepsilon)\cdot B$. At this point the high transaction cannot get scheduled, since scheduling it will just exceed the maximum allowed block size of $2\cdot B$, so the block closes at size $(1+\varepsilon)\cdot B$. Thus the next price will be $p_{t+1}=p^{t}\cdot e^{\varepsilon\eta}$. Notice that at this rate of price increase, it will take $\Omega(1/(\varepsilon\eta))$ blocks for the price to reach the value 2. As long as this is the case, the EIP-1559 algorithm’s welfare per block is less than a quarter of the optimal schedule’s welfare per block (since the optimal schedule can at the very least accept a high transaction of value 10 every block while the algorithm accepts transactions of value $2\cdot(1+\varepsilon)$ which for small enough $\varepsilon$ is less than $2.5=10/4$). Now, if we take $T$ to be half of the time it took for the price to rise to $2$ (so, $T=\Omega(1/(\varepsilon\eta))$) we must have the extension also be at least $\Gamma\geq T$ in order to reach even half of the optimal social welfare (since the algorithm that accrues welfare at a rate of a quarter of the optimal, must run at least for twice as many blocks in order to reach half the social welfare). In order to show that no extension $\Gamma(T)=o(T)$ suffices to get even half of the social welfare on this example, just choose $\varepsilon=O(1/(\eta T))$ for $T$ so that $\Gamma(T)\leq T$ [^26] When $c>2$ but is close to 2, one may take this example with $\varepsilon\approx c-2$ and get a lower bound of $\Omega\left(\frac{1}{\eta(c-2)}\right)$ on the extension.
@@
Thus, this lower bound shows that Ethereum’s choice of parameter $c=2$ is problematic when there are transactions of size $q_{max}=B$ that contributes significantly to the welfare. As the general version of our main theorem showed, this problem disappears once transaction sizes are limited by some $q_{max}<B$.
@@ 4.3.2 EIP-1559 suffers a single logarithmic loss in the range size
We show that these upper bounds are the correct ones for the EIP-1559 algorithm. Our example starts with steps where no transactions arrive, and so blocks are empty and the price decreases to $p_{min}\leq 1$. At this point of time, a large amount of “high” transactions with value per unit of $v=H$ arrive. Additionally, at that time also arrive a large amount of “low” transactions with value of $L$ per unit, where $H>L>1$ are parameters. The exact size composition of low and high transactions does not matter, as long as there is enough demand in each of these values to always completely fill blocks to the maximum block size. For our example, the low transactions will offer high “tips”, i.e. the adversary will give them priority.
So how will the EIP-1559 algorithm react to this input? After the price decreases to $p_{min}\leq 1$, only low transactions are scheduled, filling blocks to the maximum block size, so the prices will increase as $p_{t+1}=p_{t}\cdot e^{(c-1)\eta}$, where $c$ is the maximum block size multiplier. So after $k$ steps we have $p_{t+k}=p_{t}\cdot e^{k\cdot(c-1)\eta}$. As long as $L\geq p_{t}$ the adversary will only schedule low transactions, while the optimum schedule will only have high transactions, getting a ratio of $H/L$ in the social welfare. This can go on for at most $k=\Theta(\log L/((c-1)\eta))$ steps (after which $p_{t}$ that started at $p_{min}\leq 1$ can exceed $L$), which by choosing $L$ to be an arbitrary constant factor lower than $H$ (so $\log H=\Theta(\log L)$) gives us that extension $\Gamma\geq k=\Omega(\log H/((c-1)\eta))$ is needed to before we get better than the constant factor approximation $H/L$ to the welfare. Also, looking at these $k$ steps of price increase, note that all blocks are completely full, i.e. their total size is $k\cdot c\cdot B$, which implies that the slackness is $\Delta\geq k\cdot(c-1)=\Omega(\log H/\eta)$.

### v2

@@ 3.3 Proof of The Main Lemma
Since $\beta>{\color[rgb]{1,0,0}\alpha+}1$ is the first time the price is at least $\theta^{\prime}$, the price was rising from time $\beta-1$, and $p_{\beta-1}<\theta^{\prime}=\theta\cdot e^{-\eta}$. In this case $p_{\beta}$ is at most $\theta^{\prime}\cdot e^{\eta\cdot(c-1)}$. As once the price gets to $\theta^{\prime}$ it never drops below it, we have $p_{T+\Gamma+1}\geq\theta^{\prime}$. We conclude that $\frac{p_{\beta}}{p_{T+\Gamma+1}}\leq\frac{\theta^{\prime}\cdot e^{\eta\cdot(c-1)}}{\theta^{\prime}}=e^{\eta\cdot(c-1)}$ and therefore
@@ 4 Lower Bounds
Our lower bounds apply to algorithms that, like the EIP-1559 algorithm, are allowed both some slackness $\Delta$ and some extension $\Gamma$ in order to compete with the optimum schedule. As these two notions of augmentation may be traded-off against each other to some extent [^24] All our lower bounds hold even for a single horizon $T$ that is known to the algorithm in advance. In that setting, shifting between the two parameters is indeed possible. If there is a single horizon $T$ that is known by the algorithm, the extension may be completely transferred to additional slackness by immediately scheduling at the last block all transactions that were intended for the extension period. Similarly, slackness can be shifted to an extension when $T$ is known, by postponing some scheduling to the extension time., we will generally prove lower bounds on their sum.
@@ 4.1 Augmentation of Price-Based Algorithms must depend on Value Range
We will look at our algorithm running for $R=T+\Gamma$ steps, where all transactions arrive at time $1$. We will have $2^{R}+1$ possible inputs scenarios, where in input scenario number $m$ ($0\leq m\leq 2^{R}$) we have, for every integer $0\leq i\leq m$, exactly $(R+\Delta)$ transactions with value $r^{i}$ (all arriving at time 1), each of size $1$. We also specify how our adversary will respond to every price $p$ in each input scenario $m$. If $p>r^{m}$ then there are no transactions that are willing to pay this price so no transactions will be scheduled. if $p\leq r^{m}$ then the adversary will always schedule only transactions with the lowest possible value (i.e. those with value $r^{i}$ for the smallest $i$ so that $r^{i}\geq p$). Note that there are enough transactions of this value to suffice for anything that can be scheduled by the end of time $R$.
Now let us look at a price-based algorithm running for $R$ steps on one of these input scenarios. For any price $p$ that is queried by the algorithm there are at most two possible answers by the adversary on any one of these scenarios: either no transaction is scheduled or as many transactions as allowed by the block’s capacity, each of the value $r^{i}$ for the $i$ completely determined by the price, is scheduled. It follows that the complete set of possible behaviors of the algorithm on all possible scenarios can be described as a binary tree of depth $R$, containing at most $2^{R}-1$ possible prices in all and splitting the possible transactions into the $2^{R}$ leaves of the tree. Since there are $2^{R}+1$ possible scenarios, two of them must map into the same leaf and thus produce the exact same schedule. So let us assume that input scenarios $m<m^{\prime}$ get the same behavior by the algorithm, which must thus never schedule any transactions with any value strictly greater than $r^{m}$ as such do not exist in scenario $m$.
@@
Now for our choice of parameters. We let $T=\Gamma+\Delta$ which ensures that the numerator in the expression above is the constant $2$, and thus, the fraction of social welfare (with respect to the benchmark of any schedule with worst-case block size limit $1$, and no slackness or augmentation) that is achieved by the algorithm is bounded from above by $2/r$. Our choice of $T$ also ensures that $R=T+\Gamma\leq 2T=2(\Gamma+\Delta)$ and since for our choice of $r$ we only needed that $r^{2^{R}}\leq H$ we may choose $r=H^{1/2^{2T}}=H^{1/2^{2(\Gamma+\Delta)}}=H^{1/4^{(\Gamma+\Delta)}}$, which concludes the proof. ∎
@@ 4.2 Online Algorithms need to Relax the Maximum Block Size
Any online algorithm that produces schedules with maximum block size $c\cdot B$, for some $c$ with $1\leq c<2$, and average block size limit $B$ with slackness $\Delta(T)=o(T)$, even with an extension $\Gamma(T)=o(T)$, must lose at least $\delta_{0}=\Omega(2-c)$ fraction of social welfare relative to schedules with worst-case block size limit $B$ (even when the values of all transactions are in {1,2}).
@@
Case I: $G\leq T/4$. In this case the total value achieved in the first $T/2$ steps is at most $2\cdot G+1\cdot(T/2-G)\leq T/2+G\leq 3T/4$. In this case consider the case where for the next $T/2$ steps, in each step a new transaction of size 1 and value $2$ enters. In this case the best that the mechanism can do is get value $2$ for each of the remaining $T/2$ steps and the extension period $\Gamma=o(T)$ for a total social welfare of at most $7T/4+o(T)$. In comparison the optimal schedule with block size 1 would schedule all the green transactions in the first half, getting a value of $2$ every step for a total of $2T$, so the ratio of the two welfare expressions is bounded from above by $7/8+o(1)$.
Case II: $G>T/4$. In this case consider the case where for the next $T/2$ steps we get a large number of small transactions of size $\gamma$ with value $\gamma$ (i.e. value per unit of $1$) each. An optimal schedule will schedule red transactions in the first $T/2$ rounds for value of $T/2$, and all $T/2$ green transactions in the second $T/2$ rounds for value of $2\cdot T/2=T$. Additionally, all leftover capacity in the blocks containing green transactions, which have total capacity of $T/2\cdot(1-c/2-\varepsilon)$, will be filled with the small transactions, getting value of 1 per unit. So the total welfare will be $3T/2+T/2\cdot(1-c/2-\varepsilon)=T\cdot(2-c/4-\varepsilon/2)$. The algorithm is getting value of only $2\cdot G+1\cdot(T/2-G)$ in the first $T/2$ rounds, and value of only $(2+(1-(c/2+\varepsilon))\cdot(T/2-G)+1\cdot G$ in the second $T/2$ rounds. It can only get an additional $o(T)$ due to the slackness and extension. So the total value is at most $3T/2+(1-c/2-\varepsilon)\cdot(T/2-G)+o(T)<3T/2+(1-c/2-\varepsilon)\cdot T/4+o(T)=T\cdot(7/4-c/8+O(\varepsilon)+o(1))$. We conclude that the ratio of the welfare of the algorithm to the welfare of the optimum is at most $\frac{T\cdot(7/4-c/8+O(\varepsilon)+o(1))}{T\cdot(2-c/4-\varepsilon/2)}=\frac{(14-c+O(\varepsilon)+o(1))}{(16-2c-4\varepsilon)}=\frac{((16-2c)-(2-c)+O(\varepsilon)+o(1))}{(16-2c-4\varepsilon)}\leq 1-\frac{2-c}{14}+O(\varepsilon)+o(1)$ (as $2>c\geq 1$). And thus the loss is at least $\frac{2-c}{14}-O(\varepsilon)-o(1)$.
@@ 4.3.1 $c=2$ Does not Suffice for the EIP-1559 Algorithm
In Proposition 4.2 we showed that the maximum block size must be chosen to be at least $2\cdot B$ in any online algorithm that attempts to compete with offline schedules whose worst case block size is $B$ (when $q_{max}=B$), i.e., in our notation, the lower bound showed that we need $c\geq 2$. Proposition 3.1 showed that $c=2$ indeed suffices for the online algorithm that we presented. However, our main theorem for $q_{max}=B$ requires the maximum block size to be strictly greater than twice the average block size limit, $c>2$, and in fact the result also uses an extension that grows in rate that is inversely proportional to c2c-2. So the question is whether having the maximum block size $2\cdot B$ in the EIP-1559 algorithm, i.e. $c=2$, suffices for near-optimality. This question is especially interesting since the actual choice of parameters in Ethereum is exactly $c=2$, as the target block size is $B=18M$ gas and the maximum block size is defined to be $2B=36M$ gas.
@@
Here is an example showing that the EIP-1559 algorithm with $c=2$, even with an extension that grows as $\Gamma(T)=o(T)$, cannot give even half of the optimum social welfare. [^25] This example also shows that even when $c>2$, as $c$ approaches $2$, the extension length must grow inversely proportional to c2c-2. Simple tweaks in the example will show that even longer extensions cannot guarantee even a lower fraction of the welfare. For notional convenience let us assume without loss of generality that $p_{min}=1$ (otherwise just scale all the values in the rest of the proof by $p_{min}$). Our example starts with steps where no transactions arrive, and so blocks are empty and the price decreases to $p_{min}=1$. From this point on, every time step $t$, we get as input one “high” transaction of size exactly $B$ and value per unit $10$, and additional “low” transactions of total size $(1+\varepsilon)\cdot B$ and value per unit $2$, where $\varepsilon>0$ is a parameter, to be specified below, that may be as small as we wish. (The size composition of the low transactions does not matter as long as their total size is $(1+\varepsilon)\cdot B$.) In our example the low transactions offer a high tip, i.e. the adversary always schedules them before the high transaction.
@@
Now let us see how the EIP-1559 algorithm deals with this input. As long as $p_{t}\leq 2$, both the high and the low transactions can be scheduled, and since the low transactions have priority they will be scheduled first, filling the block to size $(1+\varepsilon)\cdot B$. At this point the high transaction cannot get scheduled, since scheduling it will just exceed the maximum allowed block size of $2\cdot B$, so the block closes at size $(1+\varepsilon)\cdot B$. Thus the next price will be $p_{t+1}=p^{t}\cdot e^{\varepsilon\eta}$. Notice that at this rate of price increase, it will take $\Omega(1/(\varepsilon\eta))$ blocks for the price to reach the value 2. As long as this is the case, the EIP-1559 algorithm’s welfare per block is less than a quarter of the optimal schedule’s welfare per block (since the optimal schedule can at the very least accept a high transaction of value 10 every block while the algorithm accepts transactions of value $2\cdot(1+\varepsilon)$ which for small enough $\varepsilon$ is less than $2.5=10/4$). Now, if we take $T$ to be half of the time it took for the price to rise to $2$ (so, $T=\Omega(1/(\varepsilon\eta))$) we must have the extension also be at least $\Gamma\geq T$ in order to reach even half of the optimal social welfare (since the algorithm that accrues welfare at a rate of a quarter of the optimal, must run at least for twice as many blocks in order to reach half the social welfare). In order to show that no extension $\Gamma(T)=o(T)$ suffices to get even half of the social welfare on this example, just choose $\varepsilon=O(1/(\eta T))$ for $T$ so that $\Gamma(T)\leq T$. [^26] When $c>2$ but is close to 2, one may take this example with $\varepsilon\approx c-2$ and get a lower bound of $\Omega\left(\frac{1}{\eta(c-2)}\right)$ on the extension.
@@
Thus, this lower bound shows that Ethereum’s choice of parameter $c=2$ is problematic when there are transactions of size $q_{max}=B$ that contribute significantly to the welfare. As the general version of our main theorem showed, this problem disappears once transaction sizes are limited by some $q_{max}<B$.
@@ 4.3.2 EIP-1559 suffers a single logarithmic loss in the range size
We show that these upper bounds are the correct ones for the EIP-1559 algorithm. Our example starts with steps where no transactions arrive, and so blocks are empty and the price decreases to $p_{min}\leq 1$. At this point in time, a large amount of “high” transactions with value per unit of $v=H$ arrive. Additionally, at that time also arrive a large amount of “low” transactions with value of $L$ per unit, where $H>L>1$ are parameters. The exact size composition of low and high transactions does not matter, as long as there is enough demand in each of these values to always completely fill blocks to the maximum block size. For our example, the low transactions will offer high “tips”, i.e. the adversary will give them priority.
So how will the EIP-1559 algorithm react to this input? After the price decreases to $p_{min}\leq 1$, only low transactions are scheduled, filling blocks to the maximum block size, so the prices will increase as $p_{t+1}=p_{t}\cdot e^{(c-1)\eta}$, where $c$ is the maximum block size multiplier. So after $k$ steps we have $p_{t+k}=p_{t}\cdot e^{k\cdot(c-1)\eta}$. As long as $L\geq p_{t}$ the adversary will only schedule low transactions, while the optimum schedule will only have high transactions, getting a ratio of $H/L$ in the social welfare. This can go on for at most $k=\Theta(\log L/((c-1)\eta))$ steps (after which $p_{t}$ that started at $p_{min}\leq 1$ can exceed $L$), which by choosing $L$ to be an arbitrary constant factor lower than $H$ (so $\log H=\Theta(\log L)$) gives us that extension $\Gamma\geq k=\Omega(\log H/((c-1)\eta))$ is needed before we get better than the constant factor approximation $H/L$ to the welfare. Also, looking at these $k$ steps of price increase, note that all blocks are completely full, i.e. their total size is $k\cdot c\cdot B$, which implies that the slackness is $\Delta\geq k\cdot(c-1)=\Omega(\log H/\eta)$.